# Equivariant Ehrhart theory, commutative algebra and invariant triangulations of polytopes

### v2

@@ 3.2. Constructing invariant triangulations
We continue with Example 2.15 and Example 2.25. Recall that $d>1$, $\psi:\mathbb{Z}^{d+1}\to\mathbb{Z}$, $\psi(u_{1},\ldots,u_{d},u_{d+1})=\sum_{i=1}^{d}u_{i}$, and $M^{\prime}=\psi^{-1}(1)$ is the affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(1)$. Recall that $\rho:G\to\operatorname{Aff}(M^{\prime})$ is the affine representation with $G=\operatorname{Sym}_{d}$ acting by permuting coordinates in the first $d$ coordinates, $Q=\operatorname{Conv}\left\{{e_{1},\ldots,e_{d}}\right\}\subset W$ and $P=Q\times[0,e_{d+1}]\subset W$. Let $H$ be any subgroup of $G$ and let $N$ be a positive integer. By Corollary 1.6 and Example 2.25, if $\exp(H)\nmid N$ then there does not exist an $H$-invariant triangulation of $P$ with vertices in $\frac{1}{N}M^{\prime}$. On the other hand, Theorem 1.5 implies that there exists a regular $H$-invariant triangulation of $P$ with vertices in $\frac{1}{N}M^{\prime}$ if $|H|$ divides $N$. In Question 3.10 above, we ask if there exists a regular $H$-invariant triangulation of $P$ with vertices in $\frac{1}{N}M^{\prime}$ if $\exp(H)|N$. We claim that this does exist when $H=G$. In this case, $\exp(G)=\operatorname{lcm}(m:1\leq m\leq d)$. To prove the claim it is enough to consider the case when $N=\exp(G)$. Starting with the trivial subdivision, apply successive pulling refinements by the $G$-orbits
@@
By Lemma 3.6, this produces a $G$-invariant regular triangulation of $P$ with vertices in $\frac{1}{\exp(G)}M^{\prime}$. Finally, we mention that the (necessarily regular) lattice triangulations of $P$ are in bijection with vertices of the secondary polytope of $P\cap M^{\prime}$ [DLRS10, Definition 5.1.6], which is combinatorially equivalent to the $d$-permutahedron [GKZ94, Example 7.3C]. Here the $d$-permutahedron is the convex hull of the orbit $G\cdot(1,2,\ldots,d)\subset\mathbb{R}^{d}$, where $G$ acts on $\mathbb{R}^{d}$ by permuting coordinates (cf. Example 2.26). Moreover, the combinatorial equivalence preserves the action of $G$. In particular, $G$ acts with trivial stabilizers on the $d!$ vertices of the $d$-permutahedron, and hence we have an alternative proof that there are no $H$-invariant lattice triangulations of $P$ provided $H$ is nontrivial. We next develop a criterion to guarantee the existence of a $G$-invariant lattice triangulation. Applications of this proposition will be explored in Section 3.3.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Let $\mathcal{S}$ be a $G$-invariant regular lattice polyhedral subdivision of $P$. Assume that for every face $F$ of $\mathcal{S}$ that is not a simplex there exists $u\in F\cap M$ such that $G\cdot u$ intersects every facet of $\mathcal{S}$ at most once. Then there exists a $G$-invariant regular lattice triangulation $\mathcal{S}^{\prime}$ of $P$ that refines $\mathcal{S}$. Moreover, if we can choose each $u$ above to be a vertex of $\mathcal{S}$, then we may choose $\mathcal{S}^{\prime}$ such that $\mathcal{S}$ and $\mathcal{S}^{\prime}$ have the same vertices and every simplex in $\mathcal{S}$ is a simplex in $\mathcal{S}^{\prime}$.
@@
We want to apply Lemma 3.6. We will construct a sequence $u_{1},\ldots,u_{r}\in P\cap M$ inductively for some nonnegative integer $r$. Recall from Lemma 3.6 that we consider the subcomplex $\mathcal{T}(i):=\{F\in\mathcal{S}:F\cap(\cup_{1\leq j\leq i}G\cdot u_{j})=\emptyset\}$ of $\mathcal{S}$ for all $0\leq i\leq r$. In particular, $\mathcal{T}(0)=\mathcal{S}$. Suppose that we have constructed $u_{1},\ldots,u_{i-1}\in P\cap M$ for some positive integer $i$. If all faces of $\mathcal{T}(i-1)$ are simplices then we set $r=i-1$. If not, then there exists a face $F$ in $\mathcal{T}(i-1)$ that is not a simplex. By assumption, there exists $u\in F\cap M$ such that $G\cdot u$ intersects every facet of $\mathcal{S}$ at most once. We set $u_{i}=u$. Since $\mathcal{T}(i)\neq\mathcal{T}(i-1)$, this process will terminate. Then $u_{1},\ldots,u_{r}\in P\cap M$ satisfy the hypotheses of Lemma 3.6. Lemma 3.6 then implies that $\mathcal{S}^{\prime}=\mathcal{S}_{G\cdot u_{1},\ldots,G\cdot u_{r}}$ is a $G$-invariant regular triangulation. By Remark 3.4, $\mathcal{S}^{\prime}$ is a lattice triangulation. By Remark 3.4, the vertices of $\mathcal{S}^{\prime}$ are contained in the vertices of $\mathcal{S}$ and the union of the orbits $G\cdot u_{1},\ldots,G\cdot u_{r}$. Assume that each $u_{i}$ is a vertex of $\mathcal{S}$. Then the elements of each orbit $G\cdot u_{i}$ are vertices of $\mathcal{S}$, and $\mathcal{S}$ and $\mathcal{S}^{\prime}$ have the same vertices. Finally, since applying a pulling refinement to a set of vertices of a simplex induces the trivial subdivision, every simplex in $\mathcal{S}$ is a simplex in $\mathcal{S}^{\prime}$ by Lemma 3.2. ∎
@@ 3.3. Translative actions
A translative action of a group on a poset was defined by D’alì and Delucchi in [DD21] after being introduced in a special case in [DR18]. These actions were further studied in [BD22] and [DD23]. For our purposes, we will be interested in the following special case.
@@
Let $G=\mathbb{Z}/2\mathbb{Z}=\langle\sigma\rangle$, let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Recall that $P^{\sigma}$ is the fixed locus of $\sigma$. Let $K$ be the subcomplex of $P$ consisting of the faces $F$ of $P$ such that $F\cap P^{\sigma}=\emptyset$. We claim that the induced action of $G$ on $K$ is translative. This follows since if $u$ is an vertex in $K$, then a face of $P$ containing $u$ and $\sigma\cdot u$ must also contain $\frac{u+\sigma\cdot u}{2}$ in $P^{\sigma}$ and hence does not lie in $K$.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Let $K$ be a subcomplex of $P$ and let $\mathcal{S}$ be a $G$-invariant translative regular lattice polyhedral subdivision of $K$. Then there exists a $G$-invariant translative regular triangulation $\mathcal{S}^{\prime}$ of $K$ that refines $\mathcal{S}$ such that $\mathcal{S}$ and $\mathcal{S}^{\prime}$ have the same vertices and every simplex in $\mathcal{S}$ is a simplex in $\mathcal{S}^{\prime}$.
@@
First consider the case $K=P$. Let $F$ be a face of $\mathcal{S}$ and let $u$ be any vertex of $F$. By Definition 3.13, $G\cdot u$ intersects every facet of $\mathcal{S}$ at most once. Proposition 3.12 implies that there exists a $G$-invariant regular triangulation $\mathcal{S}^{\prime}$ of $P$ that refines $\mathcal{S}$ such that $\mathcal{S}$ and $\mathcal{S}^{\prime}$ have the same vertices and every simplex in $\mathcal{S}$ is a simplex in $\mathcal{S}^{\prime}$. The action of $G$ on $\mathcal{S}^{\prime}$ is translative by Remark 3.16. Now consider the case when $K\neq P$. We claim that this reduces to the case above. As in Remark 3.1, there exists a $G$-invariant function $\omega_{K}:K\cap M\to\mathbb{R}$ such that $\mathcal{S}|_{F}=\mathcal{S}(\omega_{K}|_{A\cap F})$ for every face $F$ of $P$ contained in $K$. Let $\omega$ be the function extending $\omega_{K}$ with value zero (or any fixed constant) at all vertices of $P$ not in $K$, and set $\hat{\mathcal{S}}=\mathcal{S}(\omega)$. Then $\hat{\mathcal{S}}$ is a $G$-invariant regular lattice polyhedral subdivision of $P$ that restricts to $\mathcal{S}$ on $K$. Let $F_{1},\ldots,F_{r}$ be representatives of the $G$-orbits of the faces of $\hat{\mathcal{S}}$ not contained in $K$. After reordering, we may assume that $\dim F_{i}\geq\dim F_{i+1}$ for $1\leq i<r$. Let $x_{i}$ be an element of $M_{\mathbb{Q}}$ in the relative interior of $F_{i}$ that is fixed by the stabilizer of $F_{i}$ e.g. the average of the vertices of $F_{i}$. We apply successive pulling refinements to obtain the $G$-invariant regular polyhedral subdivision $\mathcal{T}:=\hat{\mathcal{S}}_{G\cdot x_{1},\ldots,G\cdot x_{r}}$. Then $\mathcal{T}$ is a barycentric subdivision of $\hat{\mathcal{S}}$ except that we exclude the faces of $\hat{\mathcal{S}}$ contained in $K$. Explicitly, if we choose any ordering of the elements of $G\cdot x_{i}$ and write $G\cdot x_{i}=\{y_{i,1},\ldots,y_{u,s_{i}}\}$ for some $s_{i}\in\mathbb{Z}_{>0}$, then $\mathcal{T}$ is obtained from $\hat{\mathcal{S}}$ by applying successive pulling triangulations at the points $y_{1,1},\ldots,y_{1,s_{1}},\ldots,y_{r,1},\ldots,y_{r,s_{r}}$. Moreover, by Lemma 3.5, every facet of $\mathcal{T}$ is the convex hull of a (possibly empty) face of $\mathcal{S}$ and some elements $y_{i_{1},j_{1}},\ldots,y_{i_{t},j_{t}}$ for some $i_{1}<\ldots<i_{t}$. In particular, if two distinct vertices of a facet of $\mathcal{T}$ lie in the same $G$-orbit, then the vertices must be contained in a face of $\mathcal{S}$. Since the action of $G$ on $\mathcal{S}$ is translative, we deduce that the action of $G$ on $\mathcal{T}$ is translative. Observe that the restriction of $\mathcal{T}$ to $K$ is $\mathcal{S}$. Let $M^{\prime}\subset M_{\mathbb{R}}$ be the $d$-dimensional $G$-invariant lattice generated by $M$ and all the (rational) vertices of $\mathcal{T}$. By the $K=P$ case applied to $\mathcal{T}$ and the lattice $M^{\prime}$, there exists a $G$-invariant translative regular triangulation $\mathcal{T}^{\prime}$ of $P$ that refines $\mathcal{T}$ such that $\mathcal{T}$ and $\mathcal{T}^{\prime}$ have the same vertices and every simplex in $\mathcal{T}$ is a simplex in $\mathcal{T}^{\prime}$. Then the restriction of $\mathcal{T}^{\prime}$ to $K$ is our desired triangulation refining $\mathcal{S}$.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Let $Q\subset P$ be a lattice polytope and let $K$ be the subcomplex of $P$ consisting of all faces $F$ of $P$ such that $F\cap Q=\emptyset$. Let $\mathcal{S}_{Q}$ and $\mathcal{S}_{K}$ be $G$-invariant regular lattice triangulations of $Q$ and $K$ respectively. Assume either the action of $G$ on $\mathcal{S}_{Q}$ or $\mathcal{S}_{K}$ is translative. Then there exists a $G$-invariant regular lattice triangulation $\mathcal{S}$ of $P$ such that $\mathcal{S}|_{Q}=\mathcal{S}_{Q}$ and $\mathcal{S}|_{K}=\mathcal{S}_{K}$. Moreover, the action of $G$ on both $\mathcal{S}_{Q}$ and $\mathcal{S}_{K}$ is translative if and only if we may choose $\mathcal{S}$ such that the action of $G$ on $\mathcal{S}$ is translative.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. Then $P\times[0,1]\subset M_{\mathbb{R}}\oplus\mathbb{R}$ is a $G$-invariant lattice polytope, where $G$ acts trivially on the last coordinate of $M_{\mathbb{R}}\oplus\mathbb{R}$. Let $Q=P\times\{0\}$ and $K=P\times\{1\}$. Then Remark 3.16 and Corollary 3.18 imply that the property of admitting a $G$-invariant translative regular lattice triangulation holds for $P$ if and only it holds for $P\times[0,1]$ (cf. Example 3.11).
@@
Let $G=\mathbb{Z}/2\mathbb{Z}$, let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation such that $M_{\mathbb{R}}$ has a unique $G$-fixed point. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensonal lattice polytope. That is, $P$ is a centrally symmetric polytope. We claim there exists a $G$-invariant regular lattice triangulation of $P$. Combined with Theorem 1.4, this generalizes Example 2.16 in this case. To prove the claim, let $c$ be the unique $G$-fixed point in $M_{\mathbb{R}}$ and let $\sigma$ be the generator of $G$. Fix any lattice point $u$ in $P$. The convex hull $Q$ of $u$ and $\sigma\cdot u$ is a $G$-invariant lattice simplex. Let $K$ be the subcomplex of $P$ consisting of all faces $F$ of $P$ such that $F\cap Q=\emptyset$. Since $c=\frac{u+\sigma\cdot u}{2}$ lies in $Q$, the action of $G$ on $K$ is translative by Example 3.15. The result now follows from Corollary 3.17 and Corollary 3.18.
@@
The following example shows that the answer to Question 3.21 is no if one insists that the $G$-invariant lattice triangulation is regular. In particular, the action of $G$ on the boundary of $P$ need not be translative.

### v3

@@ 3.2. Constructing invariant triangulations
We continue with Example 2.15 and Example 2.25. Recall that $d>1$, $\psi:\mathbb{Z}^{d+1}\to\mathbb{Z}$, $\psi(u_{1},\ldots,u_{d},u_{d+1})=\sum_{i=1}^{d}u_{i}$, and $M^{\prime}=\psi^{-1}(1)$ is the affine lattice in the real affine space $W=\psi_{\mathbb{R}}^{-1}(1)$. Recall that $\rho:G\to\operatorname{Aff}(M^{\prime})$ is the affine representation with $G=\operatorname{Sym}_{d}$ acting by permuting coordinates in the first $d$ coordinates, $Q=\operatorname{Conv}\left\{{e_{1},\ldots,e_{d}}\right\}\subset W$ and $P=Q\times[0,e_{d+1}]\subset W$. Let $H$ be any subgroup of $G$ and let $N$ be a positive integer. By Corollary 1.6 and Example 2.25, if $\exp(H)\nmid N$ then there does not exist an $H$-invariant triangulation of $P$ with vertices in $\frac{1}{N}M^{\prime}$. On the other hand, Theorem 1.5 implies that there exists a regular $H$-invariant triangulation of $P$ with vertices in $\frac{1}{N}M^{\prime}$ if $|H|$ divides $N$.
In Question 3.10 above, we ask if there exists a regular $H$-invariant triangulation of $P$ with vertices in $\frac{1}{N}M^{\prime}$ if $\exp(H)|N$. We claim that this does exist when $H=G$. In this case, $\exp(G)=\operatorname{lcm}(m:1\leq m\leq d)$. To prove the claim it is enough to consider the case when $N=\exp(G)$. Starting with the trivial subdivision, apply successive pulling refinements by the $G$-orbits
@@
By Lemma 3.6, this produces a $G$-invariant regular triangulation of $P$ with vertices in $\frac{1}{\exp(G)}M^{\prime}$.
Finally, we mention that the (necessarily regular) lattice triangulations of $P$ are in bijection with vertices of the secondary polytope of $P\cap M^{\prime}$ [DLRS10, Definition 5.1.6], which is combinatorially equivalent to the $d$-permutahedron [GKZ94, Example 7.3C]. Here the $d$-permutahedron is the convex hull of the orbit $G\cdot(1,2,\ldots,d)\subset\mathbb{R}^{d}$, where $G$ acts on $\mathbb{R}^{d}$ by permuting coordinates (cf. Example 2.26). Moreover, the combinatorial equivalence preserves the action of $G$. In particular, $G$ acts with trivial stabilizers on the $d!$ vertices of the $d$-permutahedron, and hence we have an alternative proof that there are no $H$-invariant lattice triangulations of $P$ provided $H$ is nontrivial.
We next develop a criterion to guarantee the existence of a $G$-invariant lattice triangulation. Applications of this proposition will be explored in Section 3.3.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Let $\mathcal{S}$ be a $G$-invariant regular lattice polyhedral subdivision of $P$. Assume that for every face $F$ of $\mathcal{S}$ that is not a simplex there exists $u\in F\cap M$ such that $G\cdot u$ intersects every facet of $\mathcal{S}$ at most once. Then there exists a $G$-invariant regular lattice triangulation $\mathcal{S}^{\prime}$ of $P$ that refines $\mathcal{S}$. Moreover, if we can choose each $u$ above to be a vertex of $\mathcal{S}$, then we may choose $\mathcal{S}^{\prime}$ such that $\mathcal{S}$ and $\mathcal{S}^{\prime}$ have the same vertices and every simplex in $\mathcal{S}$ is a simplex in $\mathcal{S}^{\prime}$.
@@
We want to apply Lemma 3.6. We will construct a sequence $u_{1},\ldots,u_{r}\in P\cap M$ inductively for some nonnegative integer $r$. Recall from Lemma 3.6 that we consider the subcomplex $\mathcal{T}(i):=\{F\in\mathcal{S}:F\cap(\cup_{1\leq j\leq i}G\cdot u_{j})=\emptyset\}$ of $\mathcal{S}$ for all $0\leq i\leq r$. In particular, $\mathcal{T}(0)=\mathcal{S}$. Suppose that we have constructed $u_{1},\ldots,u_{i-1}\in P\cap M$ for some positive integer $i$. If all faces of $\mathcal{T}(i-1)$ are simplices then we set r=i1r=i-1. If not, then there exists a face $F$ in $\mathcal{T}(i-1)$ that is not a simplex. By assumption, there exists $u\in F\cap M$ such that $G\cdot u$ intersects every facet of $\mathcal{S}$ at most once. We set $u_{i}=u$. Since $\mathcal{T}(i)\neq\mathcal{T}(i-1)$, this process will terminate. Then $u_{1},\ldots,u_{r}\in P\cap M$ satisfy the hypotheses of Lemma 3.6. Lemma 3.6 then implies that $\mathcal{S}^{\prime}=\mathcal{S}_{G\cdot u_{1},\ldots,G\cdot u_{r}}$ is a $G$-invariant regular triangulation. By Remark 3.4, $\mathcal{S}^{\prime}$ is a lattice triangulation. By Remark 3.4, the vertices of $\mathcal{S}^{\prime}$ are contained in the vertices of $\mathcal{S}$ and the union of the orbits $G\cdot u_{1},\ldots,G\cdot u_{r}$. Assume that each $u_{i}$ is a vertex of $\mathcal{S}$. Then the elements of each orbit $G\cdot u_{i}$ are vertices of $\mathcal{S}$, and $\mathcal{S}$ and $\mathcal{S}^{\prime}$ have the same vertices. Finally, since applying a pulling refinement to a set of vertices of a simplex induces the trivial subdivision, every simplex in $\mathcal{S}$ is a simplex in $\mathcal{S}^{\prime}$ by Lemma 3.2. ∎
@@ 3.3. Translative actions
A translative action of a group on a poset was defined by D’alì and Delucchi in [DD21] after being introduced in a special case in [DR18]. These actions were further studied in [BD24] and [DD23]. For our purposes, we will be interested in the following special case.
@@
Let $G=\mathbb{Z}/2\mathbb{Z}=\langle\sigma\rangle$, let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Recall that $P^{\sigma}$ is the fixed locus of $\sigma$. Let $K$ be the subcomplex of $P$ consisting of the faces $F$ of $P$ such that $F\cap P^{\sigma}=\emptyset$. We claim that the induced action of $G$ on $K$ is translative. This follows since if $u$ is an vertex in $K$, then a face of $P$ containing $u$ and $\sigma\cdot u$ must also contain $\frac{u+\sigma\cdot u}{2}$ in $P^{\sigma}$ and hence does not lie in $K$.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Let $K$ be a subcomplex of $P$ and let $\mathcal{S}$ be a $G$-invariant translative regular lattice polyhedral subdivision of $K$. Then there exists a $G$-invariant translative regular triangulation $\mathcal{S}^{\prime}$ of $K$ that refines $\mathcal{S}$ such that $\mathcal{S}$ and $\mathcal{S}^{\prime}$ have the same vertices and every simplex in $\mathcal{S}$ is a simplex in $\mathcal{S}^{\prime}$.
@@
First consider the case $K=P$. Let $F$ be a face of $\mathcal{S}$ and let $u$ be any vertex of $F$. By Definition 3.13, $G\cdot u$ intersects every facet of $\mathcal{S}$ at most once. Proposition 3.12 implies that there exists a $G$-invariant regular triangulation $\mathcal{S}^{\prime}$ of $P$ that refines $\mathcal{S}$ such that $\mathcal{S}$ and $\mathcal{S}^{\prime}$ have the same vertices and every simplex in $\mathcal{S}$ is a simplex in $\mathcal{S}^{\prime}$. The action of $G$ on $\mathcal{S}^{\prime}$ is translative by Remark 3.16.
Now consider the case when $K\neq P$. We claim that this reduces to the case above. As in Remark 3.1, there exists a $G$-invariant function $\omega_{K}:K\cap M\to\mathbb{R}$ such that $\mathcal{S}|_{F}=\mathcal{S}(\omega_{K}|_{A\cap F})$ for every face $F$ of $P$ contained in $K$. Let $\omega$ be the function extending $\omega_{K}$ with value zero (or any fixed constant) at all vertices of $P$ not in $K$, and set $\hat{\mathcal{S}}=\mathcal{S}(\omega)$. Then $\hat{\mathcal{S}}$ is a $G$-invariant regular lattice polyhedral subdivision of $P$ that restricts to $\mathcal{S}$ on $K$.
@@
Let $F_{1},\ldots,F_{r}$ be representatives of the $G$-orbits of the faces of $\hat{\mathcal{S}}$ not contained in $K$. After reordering, we may assume that $\dim F_{i}\geq\dim F_{i+1}$ for $1\leq i<r$. Let $x_{i}$ be an element of $M_{\mathbb{Q}}$ in the relative interior of $F_{i}$ that is fixed by the stabilizer of $F_{i}$ e.g. the average of the vertices of $F_{i}$. We apply successive pulling refinements to obtain the $G$-invariant regular polyhedral subdivision $\mathcal{T}:=\hat{\mathcal{S}}_{G\cdot x_{1},\ldots,G\cdot x_{r}}$. Then $\mathcal{T}$ is a barycentric subdivision of $\hat{\mathcal{S}}$ except that we exclude the faces of $\hat{\mathcal{S}}$ contained in $K$. Explicitly, if we choose any ordering of the elements of $G\cdot x_{i}$ and write $G\cdot x_{i}=\{y_{i,1},\ldots,y_{u,s_{i}}\}$ for some $s_{i}\in\mathbb{Z}_{>0}$, then $\mathcal{T}$ is obtained from $\hat{\mathcal{S}}$ by applying successive pulling triangulations at the points $y_{1,1},\ldots,y_{1,s_{1}},\ldots,y_{r,1},\ldots,y_{r,s_{r}}$. Moreover, by Lemma 3.5, every facet of $\mathcal{T}$ is the convex hull of a (possibly empty) face of $\mathcal{S}$ and some elements $y_{i_{1},j_{1}},\ldots,y_{i_{t},j_{t}}$ for some $i_{1}<\ldots<i_{t}$. In particular, if two distinct vertices of a facet of $\mathcal{T}$ lie in the same $G$-orbit, then the vertices must be contained in a face of $\mathcal{S}$. Since the action of $G$ on $\mathcal{S}$ is translative, we deduce that the action of $G$ on $\mathcal{T}$ is translative. Observe that the restriction of $\mathcal{T}$ to $K$ is $\mathcal{S}$. Let $M^{\prime}\subset M_{\mathbb{R}}$ be the $d$-dimensional $G$-invariant lattice generated by $M$ and all the (rational) vertices of $\mathcal{T}$. By the $K=P$ case applied to $\mathcal{T}$ and the lattice $M^{\prime}$, there exists a $G$-invariant translative regular triangulation $\mathcal{T}^{\prime}$ of $P$ that refines $\mathcal{T}$ such that $\mathcal{T}$ and $\mathcal{T}^{\prime}$ have the same vertices and every simplex in $\mathcal{T}$ is a simplex in $\mathcal{T}^{\prime}$. Then the restriction of $\mathcal{T}^{\prime}$ to $K$ is our desired triangulation refining $\mathcal{S}$.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Let $Q\subset P$ be a lattice polytope and let $K$ be the subcomplex of $P$ consisting of all faces $F$ of $P$ such that $F\cap Q=\emptyset$. Let $\mathcal{S}_{Q}$ and $\mathcal{S}_{K}$ be $G$-invariant regular lattice triangulations of $Q$ and $K$ respectively. Assume either the action of $G$ on $\mathcal{S}_{Q}$ or $\mathcal{S}_{K}$ is translative. Then there exists a $G$-invariant regular lattice triangulation $\mathcal{S}$ of $P$ such that $\mathcal{S}|_{Q}=\mathcal{S}_{Q}$ and $\mathcal{S}|_{K}=\mathcal{S}_{K}$. Moreover, the action of $G$ on both $\mathcal{S}_{Q}$ and $\mathcal{S}_{K}$ is translative if and only if we may choose $\mathcal{S}$ such that the action of $G$ on $\mathcal{S}$ is translative.
@@
Let $G$ be a finite group. Let $M$ be a lattice of rank $d$ and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Then $P\times[0,1]\subset M_{\mathbb{R}}\oplus\mathbb{R}$ is a $G$-invariant lattice polytope, where $G$ acts trivially on the last coordinate of $M_{\mathbb{R}}\oplus\mathbb{R}$. Let $Q=P\times\{0\}$ and $K=P\times\{1\}$. Then Remark 3.16 and Corollary 3.18 imply that the property of admitting a $G$-invariant translative regular lattice triangulation holds for $P$ if and only it holds for $P\times[0,1]$ (cf. Example 3.11).
@@
Let $G=\mathbb{Z}/2\mathbb{Z}=\langle\sigma\rangle$, let $M$ be a lattice of rank $d$, and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. Assume that the fixed locus $P^{\sigma}$ is a lattice polytope. Let $K$ be the subcomplex of $P$ consisting of the faces $F$ of $P$ such that $F\cap P^{\sigma}=\emptyset$. By Example 3.15, the induced action of $G$ on $K$ is translative. By Corollary 3.17 and Corollary 3.18, there exists a $G$-invariant regular lattice triangulation of $P$. In particular, it follows from Theorem 1.4 that $h^{*}(P,\rho;t)$ is effective.
Example 3.21.
Let $G=\mathbb{Z}/2\mathbb{Z}$, let $M$ be a lattice of rank $d$, and let $\rho:G\to\operatorname{Aff}(M)$ be an affine representation such that $M_{\mathbb{R}}$ has a unique $G$-fixed point. Let $P\subset M_{\mathbb{R}}$ be a $G$-invariant $d$-dimensional lattice polytope. That is, $P$ is a centrally symmetric polytope. We claim there exists a $G$-invariant regular lattice triangulation of $P$. Combined with Theorem 1.4, this generalizes Example 2.16 in this case.
To prove the claim, let $c$ be the unique $G$-fixed point in $M_{\mathbb{R}}$ and let $\sigma$ be the generator of $G$. Fix any lattice point $u$ in $P$. The convex hull $Q$ of $u$ and $\sigma\cdot u$ is a $G$-invariant lattice simplex. Let $K$ be the subcomplex of $P$ consisting of all faces $F$ of $P$ such that $F\cap Q=\emptyset$. Since $c=\frac{u+\sigma\cdot u}{2}$ lies in $Q$, the action of $G$ on $K$ is translative by Example 3.15. The result now follows from Corollary 3.17 and Corollary 3.18.
@@
The following example shows that the answer to Question 3.22 is no if one insists that the $G$-invariant lattice triangulation is regular. In particular, the action of $G$ on the boundary of $P$ need not be translative.